OpenStudy (anonymous):

I need help with this f(x)=2x^2+8x-8

OpenStudy (lalaly):

ok whats the question?

OpenStudy (anonymous):

graph each parabola and give its vertex,axis,x-intercepts,and y-intercept.

OpenStudy (anonymous):

x intercept set y=0 y intercept set x = 0 Vortex can be found by using the equation -(2b/a) = x where a = 2 b = +8 in this case, to find the y value sub the x value you get from that equation and you have your vortex To graph just make a table of x values to from -3 to +3 and sub them into the function to see what y values you get then graph the points and intercepts

OpenStudy (anonymous):

use quadric formula solve\[\frac{-b \pm \sqrt{b^2-(4*a*c)}}{2*a}\]a=2 b=8 c=-8

OpenStudy (anonymous):

You would use the quadratic formula to solve the x intercepts

OpenStudy (anonymous):

also whenever dealing with the quadratic formula solve what is in the radical first, dont just punch all that stuff into your calculator

OpenStudy (anonymous):

ok thanks guys

OpenStudy (anonymous):

How i can write ± in my calculator???

OpenStudy (anonymous):

u dont

OpenStudy (anonymous):

ok I will try

OpenStudy (anonymous):

solve it algebraically don't use your calculator

OpenStudy (anonymous):

Finally I got it Thanks guys